OpenStudy (anonymous):

Use properties of logarithms to expand the loarithmic expressions as much as possile. log_4 7-2 1) 7log4 4 2) 4log4 7 3) -4log4 7 4) -16log 7

OpenStudy (saifoo.khan):

7-2?

OpenStudy (anonymous):

7-^2

OpenStudy (saifoo.khan):

\[\log_4 7^{-2} \to \log_4 (\frac{1}{49})\to \frac{\log \frac{1}{49}}{\log 4}\]

OpenStudy (saifoo.khan):

are the options you wrote correct? @Audrae_World

OpenStudy (anonymous):

Yes, that's what my book says

OpenStudy (anonymous):

Isn't it \[\log_{4} 1 - \log_{4} 7^{2}\]

OpenStudy (anonymous):

Its expand not solve @saifoo.khan

OpenStudy (saifoo.khan):

Yes! im stuck. :/

OpenStudy (anonymous):

@Audrae_World : Is it \[\log_{4} 7^2\]

OpenStudy (anonymous):

That's the closest I've seen so far, but the one of answer choices says: 4log_4 7^2

OpenStudy (anonymous):

Are you sure the problem is \[\log_{4} 7^{-2}\] ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Oh. I got it. Hold on.

OpenStudy (anonymous):

\[\log_{4}7^{-2} = -2(\log_{4}7) = -2(\log_{7}7/\log_{7}4) = -2(1/\log_{7}2^{2})\] \[-2(\log_{7}2^{-2}) = -2 * -2(\log_{7}2) = 4\log_{7}2\]

OpenStudy (anonymous):

Wait. Darn it...

OpenStudy (anonymous):

LOL @Calcmathlete :whats that !??

OpenStudy (anonymous):

lol. I got so close!!! But then I realized that it isn't one of the answers...